Walter builds birdhouses and buy material $5 for each birdhouse. He builds one in 30 minutes and earns $8 per hour and sell a birdhouse for $20 each.

Calculate the total profit for one birdhouse $7,$11,$12,$15
1
Expert's answer
2011-12-08T09:02:32-0500
If we don't take into account Walter's earnings per hour the total profit for
one birdhouse will be ATP = ATR - ATC = P - ATC = 20 - 5 = $15.
If we take
into account Walter's earnings per hour the total profit for one birdhouse will
be ATP = ATR - ATC = P - ATC = 20 + 8/2 - 5 = $19.